# EN-T: Optimizing Tensor Computing Engines Performance via Encoder-Based Methodology

### v2

@@ 4.2. Performance of the Encoder
Controller and Img2col
Number Area($\mu m^{2}$) Power($W$)
2 83679 0.0632
Encoder
Number Area($\mu m^{2}$) Power($W$)
32 1895.36 0.00089
Table 2. On-chip Parameters of the SoC Benchmark
Figure 8. Benchmark SoC
4.3. Performance of EN-T architecture
Figure 9. SoC Normalized Energy Fraction under Baseline-based TCU
Figure 10. Single Frame Energy Consumption Comparison for SoC Inference Based on EN-T
[…]
@@ 4.3. Performance of EN-T architecture
4.4. Benchmark on SoC
In this section, we will evaluate the performance of the EN-T architecture on System-on-Chip (SoC) level for neural network inference (ResNet34, ResNet50, ResNet101, Inception_V3, DenseNet121, DenseNet161, Vgg13, Vgg19). We use a basic NPU architecture and run on 500MHz, as shown in the Fig.8, which includes three levels of storage. The SoC contains two levels of on-chip storage: a 256KB Global Buffer and 64KB Activation and Weight Buffers. In the readout of the Weight Buffer, we have added 32 Encoders to convert weights into encoded numbers for computation in the TCUs, with the Encoder module using register output. The Controller module includes control for reading and writing to SRAM and contains an img2col module for preprocessing the convolution operations. We use five types of TCU architectures: 2D Matrix, 1D/2D Array, two types of Systolic Array (WS and OS), and 3D Cube, with an array size of 32$\times$32. The 3D Cube configuration consists of two $8^{3}$ arrays, with a unified computational scale of 1024GOPS. Externally to the TCUs, there is a SIMD Vector Processing Engine, internally using 32 ALUs with TF32 precision for quantization, pooling, scalar addition, and activation functions. The detailed hardware parameters of each module are shown in Table.2.
Figure 11. Energy Consumption Reduction Ratio of SoC
We conducted an energy consumption analysis for single-frame image (1,3,224,224) inference on the entire SoC, starting with an analysis of the energy proportion normalized for the SoC based on the Baseline TCU, as shown in Fig.9. We decomposed the on-chip energy consumption into the read and write energy of SRAM, as well as the energy generated by the computing engines (TCU and SIMD Vector Processing Engine). For neural network computations, the energy consumed by the computing engines accounts for a high proportion of the on-chip energy, ranging from 80-94%. This is due to the high data reuse rate inherent in convolution and matrix multiplication. For instance, in a Systolic Array, each read and write operation to SRAM allows data to be transmitted and undergo multiply-accumulate operations across all rows and columns of the array, making the computing array, rather than on-chip memory access, the primary source of power consumption. However, for some lightweight networks like DenseNet or Mobilenet that use a higher proportion of memory-intensive depthwise separable convolutions, the proportion of power consumption from memory access increases, but still does not exceed 25% (as shown in Fig.9(c)). Therefore, reducing the power consumption of the TCU is beneficial for lowering the system.
@@ 4.4. Benchmark on SoC
Figure 12. Area Efficiency of the SoC
In terms of area efficiency, we compared the area efficiency improvement ratios of the Baseline as show in Fig.12, as well as the individual TCU and SoC. Due to the addition of on-chip SRAM, Controller, and SIMD Vector Processing Engine, as shown in Table.2, the area proportion of the on-chip SRAM is basically equivalent to that of the computing modules. Therefore, from the perspective of SoC, the area benefits brought by applying the EN-T architecture are relatively low. The main advantage is that it can reduce the inference power consumption by 10%-16%.
5. Conclusion
This paper explores new avenues in addressing the computational reuse issues in traditional TCUs by designing a matrix multiplication unit with an EN-T architecture. This method boasts high versatility and can be directly applied to existing tensor computing engines. From the perspective of computational encoding, we propose a method for encoding data with lower bit widths within the multiplier and uncover new opportunities for reducing energy consumption from the perspective of the multiplier’s internals. Our method combines the multiplier and adder calculation with matrix multiplication from a more fine-grained perspective and designs a low-bit width data coding method based on the carry chain. Overall, our scheme can inspire other researchers to reconsider the design of tensor computing units.

### v3

@@ III-C Performance of EN-T architecture
TABLE II: On-chip Parameters of the SoC Benchmark
Figure 9: Single frame energy consumption comparison for SoC inference based on EN-T architecture.
Figure 10: SoC normalized energy fraction under baseline-based TCU.
@@
III-D Benchmark on SoC
In this section, we will evaluate the performance of the EN-T architecture on System-on-Chip (SoC) level for neural network inference (ResNet34, ResNet50, ResNet101, Inception_V3, DenseNet121, DenseNet161, Vgg13, Vgg19). We use a basic NPU architecture and run on 500MHz, as shown in the Fig.8, which includes three levels of storage. The SoC contains two levels of on-chip storage: a 256KB Global Buffer and 64KB Activation and Weight Buffers. In the readout of the Weight Buffer, we have added 32 Encoders to convert weights into encoded numbers for computation in the TCUs, with the Encoder module using register output. The Controller module includes control for reading and writing to SRAM and contains an img2col module for preprocessing the convolution operations. We use five types of TCU architectures: 2D Matrix, 1D/2D Array, two types of Systolic Array (WS and OS), and 3D Cube, with an array size of 32$\times$32. The 3D Cube configuration consists of two $8^{3}$ arrays, with a unified computational scale of 1024GOPS. Externally to the TCUs, there is a SIMD Vector Processing Engine, internally using 32 ALUs with TF32 precision for quantization, pooling, scalar addition, and activation functions. The detailed hardware parameters of each module are shown in Table.II.
Figure 11: Energy consumption reduction ratio of SoC.
We conducted an energy consumption analysis for single-frame image (1,3,224,224) inference on the entire SoC, starting with an analysis of the energy proportion normalized for the SoC based on the baseline TCU, as shown in Fig.10. We decomposed the on-chip energy consumption into the read and write of SRAM, as well as the energy generated by the computing engines (TCU and SIMD Vector Processing Engine). For neural networks, the energy consumed by the computing engines accounts for a high proportion of the on-chip energy, ranging from 80-94%. This is due to the high data reuse rate inherent in convolution and matrix multiplication. For instance, in a Systolic Array, each read and write operation to SRAM allows data to be transmitted and undergo multiply-accumulate operations across all rows and columns of the array, making the computing array become the primary source of power consumption. However, for some lightweight networks like DenseNet or Mobilenet that use a higher proportion of memory-intensive depthwise separable convolutions, the proportion of power consumption from memory access increases, but still does not exceed 25% (as shown in Fig.10(c)). Therefore, reducing the power consumption of the TCU is beneficial for lowering the system.
@@ III-D Benchmark on SoC
In terms of area efficiency, we compared the area efficiency improvement ratios of the baseline as show in Fig.12, as well as the individual TCU and SoC. Due to the addition of on-chip SRAM, Controller, and SIMD Vector Processing Engine, as shown in Table.II, the area proportion of the on-chip SRAM is basically equivalent to that of the computing modules. Therefore, from the perspective of SoC, the area benefits brought by applying the EN-T architecture are relatively low. The main advantage is that it can reduce the inference power consumption by 10%-16%.
Figure 12: Area efficiency of the SoC.
IV Conclusion
This paper explores new avenues in addressing the computational reuse issues in traditional TCUs by designing a matrix multiplication unit with an EN-T architecture. This method boasts high versatility and can be directly applied to existing tensor computing engines. From the perspective of computational encoding, we propose a method for encoding data with lower bit widths within the multiplier and uncover new opportunities for reducing energy consumption from the perspective of the multiplier’s internals.